OpenStudy (konradzuse):

When doing a question like arctan(sec(x/8)) and doing the a^2 + b^2 = c^2

OpenStudy (konradzuse):

\[a^2 = x^2 - 8^2 -> a = \sqrt{x^2-16} or .... a = \sqrt{x^2-8^2}\]

OpenStudy (konradzuse):

I had 2 questions like this on my test on this question I did x^2-8^2 but on another one I had (2x^2)^2 and did 4x^4...... Not sure if it matters either way tho.

OpenStudy (konradzuse):

????

OpenStudy (experimentx):